Asheville Office Equipment, Inc. manufactures custom filing cabinets. Alma Ortiz, president of Asheville Office Equipment, Inc., has gathered the following cost information from the company’s accounting records for the latest month of operations.

Advertising $15,000
Fire insurance premium for the factory equipment 5,000
Sales office depreciation 2,000
Oil for the machines used to produce the cabinets 1,500
Wood used to produce the cabinets 30,000
Property taxes on the factory building 8,000
Freight to ship the cabinets to customers 2,500
Sales department salaries 34,000
Production supervisors’ salaries 18,000
Assembly line workers’ wages 45,000

a) Calculate total period costs

b) Calculate total product costs

c) Calculate manufacturing overhead incurred
Answer #1

a.

Total Period cost = Advertising + Sales office depreciation + Freight to ship the cabinets to customers + Sales department salaries

= 15,000 + 2,000 + 2,500 + 34,000

= $53,500

b.

Total product cost = Fire insurance premium for the factory equipment + Oil for the machines used to produce the cabinets + Wood used to produce the cabinets + Property taxes on the factory building + Production supervisors’ salaries + Assembly line workers’ wages

= 5,000 + 1,500 + 30,000 + 8,000 + 18,000 + 45,000

= $107,500

c.

Manufacturing overhead = Fire insurance premium for the factory equipment + Oil for the machines used to produce the cabinets + Property taxes on the factory building + Production supervisors’ salaries

= 5,000 + 1,500 + 8,000 + 30,000

= $44,500